OpenStudy (anonymous):

Find the cost to rent a car for 5 days at $27.50 per day and $ .16 per mile if the car is driven 378 miles.

OpenStudy (stamp):

algebraically represent $27.50 per day

OpenStudy (anonymous):

@stamp what do you mean?

OpenStudy (stamp):

how do you algebraically compute the cost for a charge of 27.50 per day?

OpenStudy (stamp):

i know that if i drive 60 miles per hour for 1 hour, i have driven 60*1 = 60 miles or if I have driven 60 miles per hour for 2 hours, I have driven 60*2 = 120 miles in otherwords, at 60 miles per hour, miles driven = 60*h m = 60h

OpenStudy (stamp):

how do I solve for cost c if i charge 27.50 per day

OpenStudy (anonymous):

You would take 27.50 times 5?

OpenStudy (stamp):

algebraically where is your variable

OpenStudy (anonymous):

C= cost? Im not sure what you are saying. sry.

OpenStudy (stamp):

c = 27.50d

OpenStudy (anonymous):